Hercules Exercise Equipment Co. purchased a computerized measuring device two years ago for $82,000. The equipment falls into the five-year category for MACRS depreciation and can currently be sold for $36,800. A new piece of equipment will cost $250,000. It also falls into the five-year category for MACRS depreciation. Assume the new equipment would provide the following stream of added cost savings for the next six years. Use Table 12–12. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods. Year Cash Savings 1 $ 63,000 2 53,000 3 51,000 4 49,000 5 46,000 6 35,000 The firm’s tax rate is 25 percent and the cost of capital is 8 percent. a. What is the book value of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) b. What is the tax loss on the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) c. What is the tax benefit from the sale? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) d. What is the cash inflow from the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) e. What is the net cost of the new equipment? (Include the inflow from the sale of the old equipment.) (Do not round intermediate calculations and round your answer to the nearest whole dollar.) f. Determine the depreciation schedule for the new equipment. (Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) g. Determine the depreciation schedule for the remaining years of the old equipment. (Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) h. Determine the incremental depreciation between the old and new equipment and the related tax shield benefits. (Enter the tax rate as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) i. Compute the aftertax benefits of the cost savings. (Enter the aftertax factor as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) j-1. Add the depreciation tax shield benefits and the aftertax cost savings to determine the total annual benefits. (Do not round intermediate calculations and round your answers to the nearest whole dollar.) j-2. Compute the present value of the total annual benefits. (Do not round intermediate calculations and round your answer to the nearest whole dollar.) k-1. Compare the present value of the incremental benefits (j) to the net cost of the new equipment (e). (Do not round intermediate calculations. Negative amount should be indicated by a minus sign. Round your answer to the nearest whole dollar.) k-2. Should the replacement be undertaken?
Answer #1
a DEPRECIATION OF OLD MACHINE
Cost of machine $82,000
Year from installation 1 2 3 4 5 6
A Depreciation Rate 20.00% 32.00% 19.20% 11.52% 11.52% 5.76%
B=82000*A Annual Depreciation $16,400 $26,240 $15,744 $9,446 $9,446 $4,723
C Accumulated Depreciation $16,400 $42,640 $58,384 $67,830 $77,277 $82,000
D=82000-C Book Value at end of year $65,600 $39,360 $23,616 $14,170 $4,723 $0
Book Value of the equipment now( at end of 2 years ) $39,360
Market Value of old eqipment $36,800
b Tax Loss on sale of old equipment $2,560 (39360-36800)
c Tax Rate=25%
Tax Benefit from Sale =2560*25%= $640
d Cash inflow from sale of old equipment =36800+640 $37,440
e Net Cost of new equipment:
Cost of new equipment $250,000
Cash inflow from sale of old equipment ($37,440)
Net Cost of new equipment: $212,560
f DEPRECIATION OF NEW MACHINE
Cost of new equipment $250,000
Year(from today) 1 2 3 4 5 6
E Depreciation Rate 20.00% 32.00% 19.20% 11.52% 11.52% 5.76%
F=250000*E Depreciation amount $50,000 $80,000 $48,000 $28,800 $28,800 $14,400
g DEPRECIATION OF OLD MACHINE FOR REMAINING YEARS
Year from today 1 2 3 4 5 6
G=From B above Annual Depreciation of old machine $15,744 $9,446 $9,446 $4,723 $0 $0
Year 1 2 3 4 5 6
h=F-G Incremental Depreciation $34,256 $70,554 $38,554 $24,077 $28,800 $14,400
I=h*25% Depreciation tax benefit (Incremental tax shield) $8,564 $17,638 $9,638 $6,019 $7,200 $3,600
Year 1 2 3 4 5 6
J Annual before tax cost saving $63,000 $53,000 $51,000 $49,000 $46,000 $35,000
i K=J*(1-0.25) After tax cost Savings $47,250 $39,750 $38,250 $36,750 $34,500 $26,250
Present value of Cash Flow=(Cash Flow)/((1+i)^N)
i=discount Rate =Cost of capital =8%=0.08
N=Year of Cash Flow
N Year From Today 1 2 3 4 5 6
L Depreciation tax shield $8,564 $17,638 $9,638 $6,019 $7,200 $3,600
M After tax cost Savings $47,250 $39,750 $38,250 $36,750 $34,500 $26,250
P=L+M Incremental Net Cash Flow $55,814 $57,388 $47,888 $42,769 $41,700 $29,850 SUM
H=G/(1.08^N) Present Value of Cash Inflow $51,680 $49,201 $38,015 $31,437 $28,380 $18,811 $217,524
PV Sumof Present Value of Cash inflows $217,524
I Initial Outlay(Net cost of new equipment) $212,560
NPV=PV-I Net Present value $4,964 (217524-920000)
The Machine Should be replaced
NPV is POSITIVE
Yes, the machine should be replaced